OpenStudy (toxicsugar22):

give the exact coordinates of any four points on the hyperbola (y+2)^2/6-(x+2)^2/10=1

OpenStudy (toxicsugar22):

hi ccan u help me

OpenStudy (anonymous):

That's pretty easy. Just choose any 4 points for x, and find the corresponding values of y. I'll do one for you. Let's take x=-2, plug in the equation and solve for x to get: \((y+2)^2/6-0=1\), which implies that \((y+2)^2=6\). Taking the square root of both sides gives \(y+2=\sqrt{6}\), and therefore \(y=\sqrt6-2 \). So, a point on the given hyperbola is \((-2,\sqrt6-2)\).

OpenStudy (toxicsugar22):

how about another point

OpenStudy (toxicsugar22):

how abut another point for x

OpenStudy (anonymous):

Show me what you can do with \(x=8\).

OpenStudy (toxicsugar22):

(y+2)^2/6-(8+2)^2/10=1

OpenStudy (toxicsugar22):

and then show me by step by step

OpenStudy (anonymous):

\[{(y+2)^2 \over 6}-{(8+2)^2 \over 10}=1 \implies {(y+2)^2 \over 6}-{100 \over 10}=1 \implies {(y+2)^2 \over 6}-10=1\] Now take 10 to the other side and then multiply both sides by 6. After that, take the square root and then subtract 2 from both sides. Please TRY :)

OpenStudy (anonymous):

*Take -10 to the other side with a positive sign*

OpenStudy (toxicsugar22):

so (y+2)^2/6=11

OpenStudy (toxicsugar22):

now what do i do

OpenStudy (anonymous):

I wrote all steps you need in my last comment; now multiply both sides by 6.

OpenStudy (toxicsugar22):

s0 (y+2)^2/36=66

OpenStudy (toxicsugar22):

is that rigtht so far

OpenStudy (anonymous):

Not quite. It should be, after multiplying by 6: \((y+2)^2=66\).

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

so now it will be (y+2)^2=8.1240

OpenStudy (anonymous):

Yeah, but this is not "exact". You should write it as \((y+2)^2=\sqrt{66}\).

OpenStudy (toxicsugar22):

so now what

OpenStudy (anonymous):

Sorry I meant \(y+2=\sqrt{66}\). Now take 2 to the other side with a minus sign, and you're done :D

OpenStudy (toxicsugar22):

y=sqrt66-2

OpenStudy (anonymous):

Yep.

OpenStudy (toxicsugar22):

yeah

OpenStudy (toxicsugar22):

ok and how about lets try 3 for x

OpenStudy (toxicsugar22):

can uwe do an easier number

OpenStudy (anonymous):

Go for i!! I am sure you can do it. You will get \(y=\sqrt{21}-2\). I have to go now :(

OpenStudy (anonymous):

\(3\) is good.